bill sold 50 tickets for 180$ if kid tickets are 3$ and adult tickets are 5$ how many did he sell of each
Answer by Maths68(1474)

Bill sold = 50 tickets
Bill sold Kid tickets = x
Bill sold Adult tickets =50-x
Kid ticket = $3
Adult ticket = $5
Total Sale = $180
(Bill sold Kid tickets)(price of a Kid ticket)+(Bill sold Adult tickets)(price of a Adult ticket)=Total Sale
(x)(3)+(50-x)(5)=180
3x+250-5x=180
-2x=180-250
-2x=-70
-2x/-2=-70/-2
x=35


Bill sold Kid tickets = x = 35
Bill sold Adult tickets =50-x = 50-35 = 15

Check
======
35*3+15*5=180
105+75=180
180=180